OpenStudy (anonymous):

Solve for t. Square root t+9=16 1.) 5 2.)7 3.) 25 4.) 49

OpenStudy (anonymous):

2

OpenStudy (anonymous):

t+9=16 Step 1: Subtract 9 from both sides. t+9−9=16−9 t=7

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

t=7 the answer is 7

OpenStudy (anonymous):

you just had to minus 9 from both sides

OpenStudy (anonymous):

Do you mind helping me with more?

OpenStudy (anonymous):

i dont mind

OpenStudy (anonymous):

but close this question and open a new one

OpenStudy (anonymous):

there?

OpenStudy (anonymous):

Okay I'll open a new one